Given the vectors A = 4i + 8j, B =  -8i + 7j and C = 12i – 10j. Find the magnitude and direction of the

    resultant by Component method.


1
Expert's answer
2021-05-08T11:48:52-0400

"\vec D=\vec A+\vec B+\vec C=(4-8+12)\vec i+(8+7-10)\vec j=8\vec i+5\vec j"


"|\vec D|=\sqrt{8^2+5^2}=9.43"


"\alpha=\tan^{-1}(\frac{5}{8})=32°" (relative to the positive direction of the x-axis)